PLEASE PLEASE HELP ME IF SOME ONE CAN
IF THE SUM OF THE FIRST N POSITIVE INTEGERS IS EQUAL TO N(N+1)/2,THEN THE SUM OF THE FIRST 25 POSITIVE INTEGERS IS.....OPTIONS ARE
(A)51
(B)52
(C)313
(D)20
(E)25

Answer by TimothyLamb(4379)

If:
SUM OF THE FIRST N POSITIVE INTEGERS = N(N+1)/2
Then:
SUM OF THE FIRST 25 POSITIVE INTEGERS = ?
---
N = 25
N(N + 1)/2 = ?
25(25 + 1)/2 = ?
25(26)/2 = ?
25*13 = ?
---
answer:
25*13 = 325
---
your problem statement is wrong!
none of the multiple choice answers you gave are correct
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php